# Imbalanced Data Clustering using Equilibrium K-Means

### v2

@@ Abstract
Traditional centroid-based clustering algorithms, such as hard K-means (HKM, or Lloyd’s algorithm) and fuzzy K-means (FKM, or Bezdek’s algorithm), display degraded performance when true underlying groups of data have varying sizes (i.e., imbalanced data). This paper introduces equilibrium K-means (EKM), a novel fuzzy clustering algorithm that has the robustness to imbalanced data by preventing centroids from crowding together in the center of large clusters. EKM is simple, alternating between two steps; fast, with the same time and space complexity as FKM; and scalable to large datasets. We evaluate the performance of EKM on two synthetic and ten real datasets, comparing it to other centroid-based algorithms, including HKM, FKM, maximum-entropy fuzzy clustering (MEFC), two FKM variations designed for imbalanced data, and the Gaussian mixture model. The results show that EKM performs competitively on balanced data and significantly outperforms other algorithms on imbalanced data. Deep clustering experiments on the MNIST dataset demonstrate the significance of making representation have an EKM-friendly structure when dealing with imbalanced data; In comparison to deep clustering with HKM, deep clustering with EKM obtains a more discriminative representation and a 35% improvement in clustering accuracy. Additionally, we reformulate HKM, FKM, MEFC, and EKM in a general form of gradient descent, where fuzziness is introduced differently and more simply than in Bezdek’s work, and demonstrate how the general form facilitates a uniform study of KM algorithms.
@@
K-means, fuzzy clustering, imbalance learning, deep clustering
@@ I Introduction
Imbalanced data refers to the true underlying groups of data having different sizes, which is common in datasets of medical diagnosis, fraud detection, and anomaly detection. Imbalanced data poses a challenge for learning algorithms because these algorithms tend to be biased towards the majority group [1]. While there is a considerable amount of research on supervised learning (e.g., classification) from imbalanced data [2, 3, 4], unsupervised learning has not been as thoroughly explored, because the unknown cluster sizes make the task more difficult [5]. Methods like resampling and boosting frequently used in supervised learning cannot be applied in unsupervised learning.
@@
Research on HKM and FKM has been ongoing. For example, the possibilistic FKM (PFKM) [11] was proposed to address the sensitivity of FKM to noise and outliers. FKM-$\sigma$ [12] was proposed to improve FKM’s performance on data points with uneven variations or non-spherical shapes in individual clusters. Fuzzy local information K-means (FLIKM) [13] was designed to promote FKM’s performance in image segmentation. Deep clustering, a technique that combines deep neural networks (DNNs) with HKM and FKM, was proposed to cluster high-dimensional data [14, 15, 16, 17].
Although many variations based on HKM and FKM have been developed to deal with different situations, most of them display degraded effectiveness in cases of imbalanced data. This is due to the so-called “uniform effect” that causes the clusters generated by these algorithms to have similar sizes even when input data has highly varying group sizes [18]. An illustration of the uniform effect of HKM and FKM is given in Fig. 1 where we can observe that the centroids of HKM and FKM crowd together in the large true cluster.
Figure 1: Clustering results of a highly imbalanced dataset. (a) Ground truth. The colors represent the labels of the data points. (b) Clustering by hard K-means. (c) Clustering by fuzzy K-means. (d) Clustering by the proposed equilibrium K-means. They are all two-step alternating algorithms that iteratively compute centroids.
@@ I-A Existing Efforts to Overcome Uniform Effect
To the best of our knowledge, there are two popular methods to overcome the uniform effect in the field of fuzzy systems. The first method is to introduce more weight on the data points in small clusters at each updating iteration, biasing learning towards them, like a modified FKM called cluster-size insensitive FKM (csiFKM) developed by Noordam et al. [19]. Later, Lin et al. [20] proposed a size-insensitive integrity-based FKM (siibFKM) based on csiFKM to reduce the sensitivity of csiFKM to the distance between adjacent clusters. However, weighting based on cluster size inadvertently increases the influence of outliers, making these algorithms sensitive to noise [21].
@@ I-B Our Contributions
In this paper, we propose equilibrium K-means (EKM) to address the issue of imbalanced data clustering and the aforementioned algorithms’ limitations. Our contributions can be summarized in three main aspects.
@@
Second, we develop EKM based on the first contribution. EKM belongs to the family of fuzzy clustering and membership defined in EKM has a clear physical meaning. Repulsive forces appear among centroids of EKM, successfully reducing the uniform effect by preventing centroids from crowding together in a large cluster (see Fig. 1d for an example). Therefore, EKM is effective in dealing with imbalanced data and is not sensitive to noise. Similar to HKM and FKM, EKM is a two-step alternating algorithm that iteratively computes centroids. In addition to having the same time ($O(N)$) and space complexity as FKM, EKM has a batch-learning version that can be applied to large datasets. We demonstrate the effectiveness of EKM on balanced and imbalanced data by conducting numerical experiments on two synthetic and ten real datasets from different domains. Due to space limitations, quantitative performance metrics are provided in the supplementary material. The scatter plots of partial clustering results we provide in the paper are sufficient to demonstrate the superiority of EKM on imbalanced data.
Finally, we investigate the joint learning of DNNs and EKM. We find that mapping high-dimensional data via DNNs to an EKM-friendly space can result in more discriminative low-dimensional representation than mapping to an HKM-friendly space. When tested on an imbalanced dataset derived from MNIST, joint learning of DNNs and EKM improves clustering accuracy by 35% compared to joint learning of DNNs and HKM.
@@ I-C Organization
We introduce HKM, FKM, and MEFC in Section II. In Section III, we derive their general form and the proposed EKM. The properties of EKM are studied in Section IV. We evaluate the performance of EKM on classic clustering tasks in Section V and on deep clustering in Section VI. Finally, we conclude in Section VII.
@@ II-B The Uniform Effect of K-Means
The uniform effect refers to the propensity to generate clusters of similar sizes, which is implicitly implied in the objective (1) of KM. For simplicity, kindly consider a case with two clusters (i.e., $K=2$). Minimizing the objective function (1) is equivalent to maximizing the following objective function [27]:
@@ III-E From Boltzmann Operator to A Novel Clustering Algorithm
This updating procedure can be reformulated into a two-step iteration procedure akin to FKM, that is
$\mathbf{c}_{k}^{(\tau+1)}=\frac{\sum_{n}w_{kn}^{(\tau)}\mathbf{x}_{n}}{\sum_{n}w_{kn}^{(\tau)}},$ (38)
where weights $w_{kn}^{(\tau)}$ are calculated by
@@
The time complexity of one iteration is $O(NK^{2})$. Although the sum of weights equals one, i.e.,
$\sum_{k=1}^{K}w_{kn}^{(\tau)}=1,\,\forall n,$ (40)
it is worth noting that these weights cannot be interpreted as probabilities or memberships since some weight values are negative. The membership of EKM is defined in Section III-G.
@@ III-F Physical Interpretation of Equilibrium K-Means
Boltzmann’s law tells that at the state of thermodynamic equilibrium, the probability of a particle occupying a specific energy level decreases exponentially with the increase of the energy value of that level. Hence, the objective function (34) equals the expectation of the entire system’s energy, and EKM seeks centroids to minimize this energy expectation. Due to this connection, we refer to the proposed algorithm (alternating between (38) and (39)) as equilibrium K-means.
@@ III-G Membership Defined in Equilibrium K-Means
According to the physical interpretation of EKM, the exponential term $e^{-\alpha d_{kn}}$ can be interpreted as the unnormalized probability of the $n$-th data point belonging to the $k$-th clusters. Hence, the membership of the $n$-th data point to the $k$-th cluster can be defined as
@@
Note that, although the membership formula of EKM is the same as that (10) of MEFC, the values are different because centroids are calculated using distinct formulas. We have no intention in this paper to compare membership defined in different fuzzy clustering algorithms; therefore, we only define membership in EKM and leave the further discussion to subsequent research.
@@ III-H Convergence of Smooth K-Means
It can be easily verified that when the smooth minimum function $h$ is LogSumExp (19) and p-Norm (20) in which case SKM (42) is equivalent to MEFC and FKM, respectively, the above convergence condition can be satisfied with any initial centroids. Although the Boltzmann operator (22) is not concave, it is a smooth approximation of a concave function (the minimum function). Hence, EKM also exhibits good convergence behavior in numerical experiments.
@@ IV-A Case Study
We fix the first centroid at $\mu_{1}$, and plot the reformulated objectives of HKM (WCSS (12)), FKM ($J_{\text{PN}}$ (29)), MEFC ($J_{\text{LSE}}$ (23)), and EKM ($J_{\text{B}}$ (34)) as a function of the position of the second centroid in Fig. 2. The four objective functions behave similarly on the first two balanced datasets, but it is worth noting that the last two imbalanced datasets. Fig. 2c and Fig. 2d show that the local and global minimum points of the objective functions of HKM, FKM, and MEFC are biased towards the center of the large cluster, i.e., $\mu_{1}$. In contrast, EKM does not have an obvious local minimum and its global minimum point aligns with the true cluster center, i.e., $\mu_{2}$, highlighting EKM’s superiority in handling imbalanced data.
Figure 2: Objectives as a function of the second centroid moving along the $x$-axis (the first centroid is fixed at $\mu_{1}$). The yellow, purple, green, and blue curves are reformulated objective functions of HKM, FKM, MEFC, and EKM, respectively. (a), (b), (c), and (d) present objective functions under different data distributions.
IV-B The Analysis of EKM’s Effectiveness on Imbalanced Data
@@
A simple example is given to plot the gradient values of different KM’s objectives. We fix two centroids at $-1$ and $+1$, respectively, and move a data point along the $x$-axis. In Fig. 3a we display the gradients of WCSS, $J_{\text{LSE}}$, $J_{\text{PN}}$, and $J_{\text{B}}$ with respect to the distance between the data point and the second centroid. As evident from the figure, data points on the side of the first centroid do not impact the second centroid of HKM, but they do attract the second centroids of FKM and MEFC. This finding supports the statement in [30] that FKM has a stronger uniform effect than HKM. On the other hand, data points near the first centroid have repulsive forces on the second centroid of EKM, which compensate for the attraction from other data points, effectively reducing the uniform effect. It is worth noting that data points near the second centroid of EKM have the strongest attraction forces to this centroid. Consequently, EKM’s centroids are anchored by their surrounding data points, making their position less influenced by noise and outliers.
@@ IV-C The Choice of $\alpha$
As a rule of thumb, when the dimension of the data space is less than or equal to three, setting $\alpha=1$ appears effective after normalizing the data to have a zero mean and unit variance for each dimension. As the data space dimension increases, the data-centroid distance $d_{kn}$ increases, necessitating a decrease in $\alpha$ to ensure that the exponential term $e^{-\alpha d_{kn}}$ falls within a normal range. Hence, in the case of a higher dimensional data clustering, we suggest initially setting $\alpha$ to ten times the data variance and gradually reducing it until a sudden increase in centroid-centroid distance is observed. Because the increase in centroid distance implies the emergence of repulsive forces that reduce the uniform effect. This strategy is applied in deep clustering on the MNIST dataset in Section VI. An adaptive selection formula of $\alpha$ and its validation are given in the supplemental material.
@@ V Numerical Experiments
Numerical experiments are conducted to compare the performance of our proposed EKM algorithm with six other centroid-based algorithms: (1) HKM, (2) FKM, (3) MEFC, (4) csiFKM [19], (5) siibFKM [20], and (6) the GMM. Multiprototype-based algorithms such as [22, 5] are not appropriate as baseline models because they are too complex to be benchmarks for gauging the efficiency of the proposed EKM.
To avoid local optima convergence, each clustering algorithm is executed ten times, and the result with the lowest objective value is presented. Centroids are initialized independently for each replication by the K-means++ algorithm [32]. Convergence is achieved when the relative moving distance of centroids between successive iterations is less than $1\mathrm{e}{-3}$, i.e.,

### v3

@@ Abstract
Centroid-based clustering algorithms, such as hard K-means (HKM) and fuzzy K-means (FKM), have suffered from learning bias towards large clusters. Their centroids tend to be crowded in large clusters, compromising performance when the true underlying data groups vary in size (i.e., imbalanced data). To address this, we propose a new clustering objective function based on the Boltzmann operator, which introduces a novel centroid repulsion mechanism, where data points surrounding the centroids repel other centroids. Larger clusters repel more, effectively mitigating the issue of large cluster learning bias. The proposed new algorithm, called equilibrium K-means (EKM), is simple, alternating between two steps; resource-saving, with the same time and space complexity as FKM; and scalable to large datasets via batch learning. We substantially evaluate the performance of EKM on synthetic and real-world datasets. The results show that EKM performs competitively on balanced data and significantly outperforms benchmark algorithms on imbalanced data. Deep clustering experiments demonstrate that EKM is a better alternative to HKM and FKM on imbalanced data as more discriminative representation can be obtained. Additionally, we reformulate HKM, FKM, and EKM in a general form of gradient descent and demonstrate how this general form facilitates a uniform study of K-means algorithms.
@@
K-means, fuzzy clustering, imbalance learning, uniform effect, deep clustering
@@ I Introduction
Imbalanced data refers to the true underlying groups of data having different sizes, which is common in datasets of medical diagnosis, fraud detection, and anomaly detection. Imbalanced data poses a challenge for learning algorithms because these algorithms tend to be biased towards the majority group [1]. While there is a considerable amount of research on supervised learning (e.g., classification) from imbalanced data [2, 3, 4], unsupervised learning has not been as thoroughly explored, because the unknown cluster sizes make the task more difficult [5]. Methods like resampling and boosting frequently used in supervised learning cannot be applied in unsupervised learning due to the lack of labels.
@@
HKM and FKM stimulate many subsequent research. For example, the possibilistic K-means (PKM) [11] was proposed with reformulated membership more meaningful in terms of typicality. Later, the possibilistic fuzzy K-means (PFKM) [12] was proposed to address the noise sensitivity defect of FKM and overcome the coincident clusters problem of PKM. FKM-$\sigma$ [13] was proposed to improve FKM’s performance on data points with uneven variations or non-spherical shapes in individual clusters. Fuzzy local information K-means [14] was designed to promote FKM’s performance in image segmentation. Recently, feature-weighted PKM (FWPKM) [15, 16] was proposed to give non-uniform importance to features. Research of combining K-means and kernel mechanisms [17, 18, 19] has garnered interest, which non-linearly mapped data from low-dimensional space to high-dimensional space through kernel functions to enhance data separability. Conversely, deep clustering, a modern technique that integrates deep neural networks (DNNs) and K-means (or other clustering algorithms), has been proposed to cluster high-dimensional data by mapping them to a low-dimensional space to overcome the curse of dimensionality [20, 21, 22, 23].
Although many variations derived from HKM and FKM have been developed to deal with different situations, most display degraded effectiveness in cases of imbalanced data. This is due to the so-called “uniform effect” that causes the clusters generated by these algorithms to have similar sizes even when input data has highly varying group sizes [24]. It is reported that FKM has a stronger uniform effect than HKM [25]. An illustration of the uniform effect of HKM and FKM is given in Fig. 1 where we can observe that the centroids of HKM and FKM crowd together in the large cluster.
Figure 1: Clustering results of a highly imbalanced dataset. (a) Ground truth. The colors represent the reference labels of the data points. (b) Clustering by hard K-means. (c) Clustering by fuzzy K-means. (d) Clustering by the proposed equilibrium K-means. They are all two-step alternating algorithms that iteratively update centroids.
@@ I-A Existing Efforts to Overcome Uniform Effect
There are two popular methods to overcome the uniform effect. The first method is to introduce more weight on the data points in small clusters at each updating iteration, biasing learning towards them, like a modified FKM called cluster-size insensitive FKM (csiFKM) developed by Noordam et al. [26]. Later, Lin et al. [27] proposed a size-insensitive integrity-based FKM (siibFKM) based on csiFKM to reduce the sensitivity of csiFKM to the distance between adjacent clusters. However, weighting based on cluster size inadvertently increases the influence of outliers, making these algorithms sensitive to noise [28].
@@ I-B Our Contributions
In this paper, we propose equilibrium K-means (EKM), which addresses the imbalanced data clustering issue by a new centroid repulsion mechanism without the above limitations. Our contributions can be summarized in three main aspects.
@@
Second, we develop EKM based on the first contribution where the Boltzmann operator is utilized as an approximation method. Such approximation brings a brand-new centroid repulsion mechanism. Data surrounding centroids repel other centroids, and large clusters repel more as they contain more data, successfully reducing the uniform effect by preventing centroids from crowding together in a large cluster (see Fig. 1d for an example). Hence, EKM is robust to imbalanced data and not sensitive to noise. Similar to HKM and FKM, EKM is a two-step alternating algorithm that iteratively computes centroids. In addition to having the same time ($O(N)$) and space complexity as FKM, EKM has a batch-learning version that can be applied to large datasets. A comprehensive study is conducted using four synthetic and 16 real datasets (hence 20 datasets in total) to demonstrate the effectiveness of EKM on balanced and imbalanced data.
Finally, we investigate the combination of DNNs and EKM. We find that mapping high-dimensional data via DNNs to an EKM-friendly space can result in more discriminative low-dimensional representation than mapping to an HKM-friendly space. Compared to the combination of DNNs and HKM, using EKM improves clustering accuracy by 35% on an imbalanced dataset derived from MNIST.
@@ I-C Organization
We introduce HKM, FKM, and MEFC in Section II. In Section III, we derive their general form and the proposed EKM. The properties of EKM and its centroid repulsion mechanism are studied in Section IV. We evaluate the performance of EKM on classic clustering tasks in Section V and on deep clustering in Section VI. Finally, we conclude in Section VII.
@@ II-B The Uniform Effect of K-Means
The uniform effect refers to the propensity to generate clusters of similar sizes. It is essentially a learning bias towards large clusters and is implicitly implied in the objective (1) of KM. For simplicity, kindly consider a case with two clusters (i.e., $K=2$). Minimizing the objective function (1) is equivalent to maximizing the following objective function [34]:
@@ III-E From Boltzmann Operator to A Novel Clustering Algorithm
This updating procedure which we call EKM can be reformulated into a two-step iteration procedure akin to FKM:
  1. 1.
    Calculate the weight value of the $n$-th data point to the $k$-th cluster by
@@
  2. 2.
    Recalculate the weighted centroid of the $k$-th cluster by
    $\mathbf{c}_{k}^{(\tau+1)}=\frac{\sum_{n}w_{kn}^{(\tau)}\mathbf{x}_{n}}{\sum_{n}w_{kn}^{(\tau)}}.$ (39)
EKM converges when centroids cease to change or the maximum number of iterations is reached. The time complexity of one iteration of the above two steps is $O(NK^{2})$. For convenience, we summarise the complete procedure of EKM in Algorithm 1.
Input: A dataset $X=\{\mathbf{x}_{n}\}_{n=1}^{N}$, cluster number $K$, initial centroids $\{\mathbf{c}_{k}^{(0)}\}_{k=1}^{K}$, smoothing parameter $\alpha$
Output: Centroids $\{\mathbf{c}_{1},\cdots,\mathbf{c}_{K}\}$
$\tau=0$;
repeat
       Compute weight $w_{kn}^{(\tau)}$ by (38) for all $k,\,n$;
       Update centroid $\mathbf{c}_{k}^{(\tau+1)}$ by (39) for all $k$;
       $\tau=\tau+1$;
until convergence;
return $\{\mathbf{c}_{k}^{(\tau)}\}_{k=1}^{K}$
Algorithm 1 Equilibrium K-Means Algorithm
@@ III-F Physical Interpretation of Equilibrium K-Means
Boltzmann’s law tells that at the state of thermodynamic equilibrium, the probability of a particle occupying a specific energy level decreases exponentially with the increase of the energy value of that level. Hence, the objective function (34) equals the expectation of the entire system’s energy, and EKM seeks centroids to minimize this energy expectation. Due to this connection, we refer to the proposed Algorithm 1 as equilibrium K-means.
@@ III-G Membership Defined in Equilibrium K-Means
Membership can be defined in EKM. Although the sum of weights (38) of EKM equals one, i.e.,
$\sum_{k=1}^{K}w_{kn}^{(\tau)}=1,\,\forall n,$ (40)
it is worth noting that these weights cannot be interpreted as probabilities or memberships since some weight values are negative. We define membership of EKM from its physical perspective. From the physical interpretation of EKM, the exponential term $e^{-\alpha d_{kn}}$ can be interpreted as the unnormalized probability of the $n$-th data point belonging to the $k$-th clusters. Hence, the membership of the $n$-th data point to the $k$-th cluster can be defined as
@@
Since we have no intention in this paper to discuss the membership (it is just an intermediate product), we only define the membership in EKM and leave further discussion to subsequent research.
@@ III-H Convergence of Smooth K-Means
It can be easily verified that when the smooth minimum function $h$ is LogSumExp (19) and p-Norm (20) in which case SKM (42) is equivalent to MEFC and FKM, respectively, the above convergence condition can be satisfied with any initial centroids. Since the Boltzmann operator (22) is not globally concave, discussing the convergence of EKM is not simple. If centroids are updated within the concave region of the Boltzmann operator, the convergence of EKM can be guaranteed by Theorem 1. However, the practical situation may be more complex. A rigorous convergence condition of EKM may involve a tedious discussion of a specific value of $\alpha$, positions of initial centroids, and data structures. Instead, we demonstrate the empirical convergence of EKM in Section V-F.
@@ IV-A Case Study
We plot $J(\mathbf{c}_{1}=\mu_{1},\mathbf{c}_{2})$ as a function of $\mathbf{c}_{2}$ in Fig. 2, with $\mathbf{c}_{2}$ being the position of the second centroid on the $x$-axix, and $J$ being the reformulated objectives of HKM (WCSS (12)), FKM ($J_{\text{PN}}$ (29)), MEFC ($J_{\text{LSE}}$ (23)), and EKM ($J_{\text{B}}$ (34)). The four objective functions behave similarly on the first two balanced datasets, but it is worth noting that the last two imbalanced datasets. Fig. 2c and Fig. 2d show that the local and global minimum points of the objective functions of HKM, FKM, and MEFC are biased towards the center of the large cluster, i.e., $\mu_{1}$. In contrast, EKM does not have an obvious local minimum and its global minimum point aligns with the true cluster center, i.e., $\mu_{2}$, highlighting EKM’s superiority in handling imbalanced data.
Figure 2: Objectives $J({\mathbf{c}_{1}=\mu_{1},\mathbf{c}_{2}})$ as a function of $\mathbf{c}_{2}$ (the position of the second centroid on the $x$-axis). The yellow, purple, green, and blue curves are reformulated objective functions of HKM, FKM, MEFC, and EKM, respectively. (a), (b), (c), and (d) present objective functions under different data distributions.
IV-B The Analysis of EKM’s Effectiveness on Imbalanced Data: A Centroid Repulsion Mechanism
@@
A simple example is given to plot the gradient values of different KM’s objectives. We fix two centroids at $-1$ and $+1$, respectively, and move a data point along the $x$-axis. In Fig. 3a we display the gradients of WCSS, $J_{\text{LSE}}$, $J_{\text{PN}}$, and $J_{\text{B}}$ with respect to the distance between the data point and the second centroid. As evident from the figure, data points on the side of the first centroid do not impact the second centroid of HKM, but they do attract the second centroids of FKM and MEFC. This finding supports the claim in [25] that FKM has a stronger uniform effect than HKM. On the other hand, data points near the first centroid have repulsive forces on the second centroid of EKM, which compensate for the attraction from other data points. This new mechanism effectively reduces the algorithm’s learning bias towards other clusters, particularly large ones, because the centers of large clusters often contain more data. The existence of this mechanism suggests that the uniform effect of EKM is weak, a hypothesis that our experiments will test. Notably, data points near the second centroid of EKM exert the strongest attraction forces on this centroid. As a result, EKM’s centroids are stabilized by their surrounding data points, reducing their susceptibility to noise and outliers.
@@ IV-C The Choice of EKM’s parameter
As a rule of thumb, when the dimension of the data space is less than or equal to three, setting $\alpha=1$ appears effective after normalizing the data to have a zero mean and unit variance for each dimension. As the data space dimension increases, the data-centroid distance $d_{kn}$ increases, necessitating a decrease in $\alpha$ to ensure that the exponential term $e^{-\alpha d_{kn}}$ falls within a normal range. Hence, one can set $\alpha$ as a reciprocal of the data variance or manually tune it as follows. Initially setting $\alpha$ as ten times the reciprocal of data variance and gradually reducing it until a sudden increase in centroid-centroid distance is observed. Because the increase in centroid distance implies the emergence of repulsive forces that reduce the uniform effect.
@@ V Numerical Experiments
V-A Experimental Setup
Numerical experiments are conducted to compare the performance of our proposed EKM algorithm with seven related centroid-based algorithms including (1) HKM, (2) FKM, (3) MEFC, (4) PFKM [12], (5) csiFKM [26], (6) siibFKM [27], and (7) FWPKM [15]. Multiprototype clustering algorithms (e.g., [29, 5]) are not appropriate as baseline algorithms because they are too complex to be benchmarks for gauging the efficiency of EKM. We also avoid using hybrid methods, such as algorithms combining K-means with kernels proposed in [17, 18, 19], as benchmarks to ensure fairness. Implementation is conducted in Matlab R2022a and the operation system is Ubuntu 18.04.1 LTS with Intel Core i9-9900K CPU @ 3.60GHz X 16 and 62.7GiB memory. All datasets used and codes are available at https://github.com/ydcnanhe/Imbalanced-Data-Clustering-using-Equilibrium-K-Means.git.
The experimental datasets contain four artificial datasets generated by us (including Data-A, Data-B, Data-C, and Data-D), 13 UCI [38] datasets (consisting of Image Segmentation (IS), Seeds, Wine, Rice, Wisconsin Diagnostic Breast Cancer (WDBC), Ecoli, Htru2, Zoo, Glass, Shill Bidding, Anuran Calls, Occupancy Detection, Machine Failure), and three Kaggle datasets (incorporating Heart Disease, Pulsar Cleaned, and Bert-Embedded Spam). So there are 20 datasets and Table I provides their information, including name, instance number, feature number, reference class number, and coefficient of variation (CV). CV is used in previous literature [34] to measure the level of data dispersion. It is calculated as the ratio of the standard deviation of class sizes to the mean. Given the number of instances in each class as $N_{1},\cdots,N_{K}$, we have
$\text{CV}=s/\bar{N},$ (44)
where
$\bar{N}=\frac{\sum_{k=1}^{K}N_{k}}{K},\,s=\sqrt{\frac{\sum_{k=1}^{K}(N_{k}-\bar{N})^{2}}{K-1}}.$
In [34], a CV value exceeding 1 indicates highly varying class sizes, while a value below 0.3 signifies uniform class sizes. However, there isn’t a widely accepted critical CV value implying that data is imbalanced. For rigorous statements, we establish that if a dataset’s CV is less than 0.4, it is considered balanced. Conversely, if the CV exceeds 0.7, the dataset is deemed imbalanced. Consequently, we have six balanced datasets and 14 imbalanced datasets. It should be noted that none of the datasets used has a CV value between 0.4 and 0.7, hence this range remains undefined.
TABLE I: Detailed description of 20 Datasets
ID Name Instances Feature Classes CV
D1 Data-A 2250 2 3 1.4468
D2 Data-B 2250 2 3 1.4468
D3 Data-C 5200 2 2 1.3054
D4 Data-D 5400 2 9 2.7500
D5 IS 2310 19 7 0
D6 Seeds 210 7 3 0
D7 Heart Disease 1125 13 2 0.0373
D8 Wine 178 13 3 0.1939
D9 Rice 3810 7 2 0.2042
D10 WDBC 569 30 3 0.3604
D11 Zoo 101 16 7 0.8937
D12 Glass 214 9 6 1.0767
D13 Ecoli 336 7 8 1.1604
D14 Htru2 17898 8 2 1.1552
D15 Shill Bidding 6321 9 2 1.1122
D16 Anuran Calls 7195 22 10 1.6016
D17 Occupancy Detection 20560 5 2 0.7608
D18 Machine Failure 9815 7 2 1.3318
D19 Pulsar Cleaned 14987 7 2 1.3561
D20 Bert-Embedded Spam 5572 768 2 1.0350
All datasets undergo normalization, ensuring that each feature has zero mean and unit variance. All features are used for clustering purposes. The number of clusters is set to the reference class number. Convergence is achieved when the moving distance of centroids between successive iterations is sufficiently small relative to the magnitude of centroids, i.e.,